On January 1, 2017, we purchased a truck for $60,000. The truck has a useful life of 5 years and a residual value of $5,000. The truck is depreciated using the straight-line method. What is the book value after the adjusting entry for 2018 (Year 2)?

a) $38,000

b) $36,000

c)$33,000

d)$32,000

On January 1, 2017, we purchased a truck for $60,000. The truck has a useful life of 5 years and a residual value of $5,000. The truck is depreciated using the double-declining balance method. What is the balance in accumulated depreciation after the adjusting entry for 2017?

a)$24,000

b)$22,000

c)$12,000

d)$11,000

A truck cost $60,000 when purchased and has a $55,000 balance in the accumulated depreciation account. If the truck is discarded we will record:

a) loss on disposal, $5,000

b) gain on disposal, $5,000

c) no gain or loss because it was discarded

d) loss of $60,000

A truck cost $60,000 when purchased and has a $55,000 balance in the accumulated depreciation account. If the truck is sold for $5,000 cash we will record:

a) gain on disposal, $5,000

b) loss on disposal, $5,000

c) no gain or loss

d) no entry required
Answer #1

1) Depreciation under straight line method = (60000-5000)/5 = 11,000 per year

Therefore depreciation for 2 years will be 22,000

Hence valuation of truck after 2 years will be 60,000 - 22,000 = 38,000

Correct answer is A

2) Under Double declining method depreciation is calculated twice the depreciation as per straight line method

Hence accumulated depreciation after 2017 will 11,000*2 = 22,000.

Correct answer is B

3) As the truck is discarded and it is not fully depreciated , therefore the difference will be net loss

Hence it will be loss of 5,000

Correct answer is A

4) As the truck is sold for 5,000 cash , therefore there will be no gain or loss

Hence the correct answer is C.

If you find this answer helpful please upvote.